OpenStudy (anonymous):

I just went brain dead and need some help. The answer keeps slipping my mind. When using a quadratic equation in the form y = ax2 + bx + c to model the height of a projectile (y) over time (x), which of the following is always represented by the constant term? the initial height of the projectile the initial velocity of the projectile the time at which the projectile hits the ground the maximum height of the projectile

Directrix (directrix):

If x represents time, at time x = 0, and y = ax2 + bx + c, then y = 0 + 0 + c = c. At time zero, the projectile has not moved, so what do you think c might represent? @GodGirl360

OpenStudy (anonymous):

The initial height

Directrix (directrix):

That is what I think.

OpenStudy (anonymous):

Ok, thank you! :)

Directrix (directrix):

You're welcome.